OpenStudy (anonymous):

Can some one PLEASE help me find all the real square roots of -9/16?

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

is it 3/4 and -3/4 or just 3/4?

OpenStudy (anonymous):

just 3/4

OpenStudy (anonymous):

but it's negative?

OpenStudy (anonymous):

i is the imaginary number defined as the square root of -1 --> i = sqrt(-1) = (-1)^1/2 knowing this, you can change sqrt(-9/16) to sqrt(-1*9/16) which equals sqrt(-1)*sqrt(9)/sqrt(16) = i*3/4 = 3/4i